OpenStudy (anonymous):

Please help, & explain your answer! Medal will be given. :) Solve 4(x + 1) = 4x + 10. 0 6 All real numbers No solution

OpenStudy (anonymous):

4(x+1)=4x+10

OpenStudy (anonymous):

from there, can you distribute the 4?

OpenStudy (anonymous):

4x + 5 = 4x + 10, correct?

OpenStudy (anonymous):

then, subtract 5 from both sides, then you have 4x = 4x + 10, right?

OpenStudy (gorv):

1*4=4 lol not 5

OpenStudy (gorv):

4x+4=4x+10 now u do it...u were goin right

OpenStudy (anonymous):

no solution bc the 4x cancels on both sides and 4 does not equal to 10

OpenStudy (gorv):

he will get there @lee989 ...

OpenStudy (anonymous):

i'm a she. :) lol thanks for the help guys, sorry i didn't respond earlier i was helping someone with a whole test. thank you! @gorv @lee989

OpenStudy (gorv):

sorry ...:P

OpenStudy (gorv):

i will medal lee than

OpenStudy (anonymous):

:)